the perimeter of a rectangle is 50 m. The width of the rectangle is 5 m less than the length. find the length and width of the rectangle
Answer by Alan3354(69443)

P = 2L + 2W = 50
L + W = 25
W = L - 5
L + (L - 5) = 25
2L -5 = 25
2L = 30
L = 15
W = 10